# Hierarchical self-assembly for high-yield addressable complexity at fixed conditions

### v1

@@ II.4 Assembling $n$-squares with high yields via hierarchical interactions
As a check that we can extend hierarchical assembly by one more generation we simulated $N_{c}=2$ copies of a 1024-square using $\delta_{1}{=}10$, for a time of $1.15\times 10^{6}$ MC sweeps, and repeated 4 times. Out of the 8 total copies of a 1024-square simulated, 7 copies formed perfectly, and 1 copy was missing two monomers from its interior. Snapshots of a perfectly-assembling system are shown in S.I., Fig. 13.
Hierarchical assembly as the number of generations $k$ varies is compared in Fig.6. To choose the duration of these runs systematically, we consider $k=1$ and determine a time $t_{1}=64\times 10^{3}$ MC sweeps at which $y_{1}$ is close enough to equilibrium over a range of $\delta_{1}$ values. For larger $k$, we chose $t_{k}=8^{k-1}t_{1}$, based on scaling time with the diffusion timescale for each generation, $\tau_{D}^{k}=(D_{k}\rho)^{-1}$ (for each subsequent generation, $\rho$ decreases by a factor of 4 for fixed volume fraction $\phi$, and $D_{k}$ decreases by a factor of 2). Generation $k=5$ is run for about half that time because of computational limitations.
Fig. 6 shows that the yields are high over a range of $\delta_{1}$-values for all generations, though they decrease slightly with each new generation as explained below. An exception is $k=5$ which has abnormally high yields, which we attribute to finite-size effects (Fig. 14; SI Section C.2). The curves appear shifted to the right by about $1k_{B}T$ for each additional generation. This is consistent with (1), which shows that for fixed error rate $\Delta_{1}$ in the first generation and fixed volume fraction $\phi=4^{k}\rho$, the interaction energy must increase by about $3/4\log 4\approx 1.04k_{B}T$ for each additional generation (see also SI, (15)).
@@ II.5 Dynamics of hierarchical assembly
Figure 9: (Top) Fraction of the system mass in fragments of size $n=4,16,64,256$, for a system of $N_{c}{=}16$ 256-squares with $\delta_{1}=8.75$, repeated 8 times. (Bottom) Same data but with time rescaled by $1/\sqrt{n}$.
@@
To quantify this behaviour further, in the top panel of Fig.9 the mass in fragments of size $n=4,16,64,256$ is shown as a function of time, for a system of assembling 256-squares with optimal hierarchical interactions. The peaks in the masses occur at subsequently increasing times and then decay, except for the targeted structure $n=256$. Noticeably, the decaying parts of the curves appear simply shifted in time by $1/\sqrt{n}$ as $n$ increases, consistent with scaling the decay with the diffusion timescales as we rationalize below.
@@ III.1 Themodynamics constraints
We seek to estimate the yield $y_{k}$ of hierarchical self-assembly with $k$ generations, as a function of the monomer density $\rho$ and of the interaction strengths. Let us denote by $\tilde{\delta}_{j}=2^{j}\delta_{j}$ the effective interaction between two fully formed squares at the $j^{th}$ intermediate generation, each consisting of $4^{j}$ monomers. We will assume that $\tilde{\delta}_{j}=\delta$ independently of $j$. We denote by $y_{j}^{(k)}$ the yield of such squares. At the first generation, assuming that weaker interactions do not disturb the kinetic of the assembly process (see below), from Eq.1 we get $y_{1}^{(k)}=1-\Delta_{1}(\rho)$. We can now proceed recursively, assuming that (i) the formation of squares at generation $j$ is not affected by that of larger structures and (ii) that it only depends on the concentration of $j-1$ squares obtained in the process, which only depends on the yields of previous stages $j^{\prime}$ with $1\leq j^{\prime}\leq j$. The yields of intermediate generations are thus approximated as
@@
and the target yield $y_{k}=y_{k}^{(k)}$ is computed recursively. This yield approximation is tested against simulation data in Fig.6 (dashed lines). It correctly captures the trend for the yield, and even agrees quantitatively for weak interactions and small $k$.
@@
Figure 10: A schematic phase diagram for hierarchical assembly.
@@ III.2 Kinetic constraints
Remaining hierarchical: For hierarchical assembly to proceed, interactions needed to build the generation $j+1$ should not interfere with the assembly of generation $j$; they should start to play a role only when the $j$-squares are fully formed. Uncompleted substructures (such as triangles) of two adjacent $j$-squares bind with an interaction $\delta/2$. They will thus stick together even when the $j$ squares are not completed, and remain attached for some time $\sim e^{\delta/2}/D_{j}$. If this time is much larger than the diffusion time $\tau_{D}^{j}\sim 1/(\rho D_{j})$ for $j$-squares to meet, then partially formed structures spend most of their time bound to other partially formed structures. Proper hierarchical assembly can only happen when this does not occur, implying:
@@
Avoiding defects: As discussed for the formation of 4-squares, the concentration of triangles will be abundant at any level $j$ at some point during the hierarchical assembly if the ratio of unbinding time over diffusion time $f^{j}\equiv\tau^{j}_{B}/\tau^{j}_{D}=e^{\delta}\rho$ is large. If the associated condition $f^{j}\gg 1$ or equivalently:
$e^{\delta}\gg\frac{1}{\rho}$ (8)
holds, we expect to see defects composed of a triangle binding to three completed squares, such that an inaccessible hole is present in the interior of this structure. Indeed, the energy of removing the triangle to heal such a defect is precisely $\delta$; condition Eq.8 ensures that the healing time is smaller than the characteristic time where intermediate structures meet and bind.
Interestingly, if we denote by $t_{j}$ the time scale to form generation $j$, we have that $t_{j}$ is always inversely proportional to $D_{j}$. Indeed, $t_{j}$ corresponds to the slowest time between the unbinding time $\tau^{j}_{B}$ and the diffusion time $\tau^{j}$, and both satisfy this property. Since $D_{j}\sim 1/\sqrt{4^{j}}$, one expects that during the hierarchical assembly, subsequent generations should take 2 times longer to form. Thus, $t_{j}\sim 2^{j}\sim\sqrt{n}$, as confirmed in Fig.9.
Figure 11: Snapshots showing qualitatively how hierarchical assembly behaves as $\delta_{1}$ increases, for $n=64$ at fixed volume fraction $\phi$. This corresponds to moving upwards along a vertical line in the phase diagram in Fig. 10. The examples are taken from simulations described in Fig. 12, and are plotted at $t=256$ (time units are described in the caption to Fig. 12).
Figure 12: Detailed simulations of 64-square assembly. Left: yield of perfectly-formed 64-squares at different times and different $\delta_{1}$-values from simulations (markers), and a spline curve fit through each set of data to show the trend (solid curves). Left: same, but the yield includes structures with at most 2 monomers missing (markers), with spline curves (dashed lines). For each value of $\delta_{1}$, simulations considered 60 copies of a 64-square, repeated 3 times, at volume fraction $\rho=0.05$. Time is measured in units of $32\times 10^{3}$ MC sweeps.
@@ III.3 Phase diagram
Our results are summarized in the phase diagram of Fig. 10, with snapshots from different regions of the diagram shown in Fig. 11. At fixed $\rho$, at low $\delta$ the yield is poor as expected thermodynamically, and improves continuously as $\delta$ increases. For $e^{\delta}>C/\rho$ (here $C$ represents an unknown constant), defects can start to appear. For even stronger interactions $e^{\delta}>C/\rho^{2}$, the notion of hierarchical assembly is lost, as interactions associated with different levels can bind concomitantly. Finally for $e^{\delta}>C\,tD$, low-level bonds have no time to detach: kinetic traps appear already at small scales.
@@
Overall, we expect that yield is governed by a trade-off between thermodynamics, which favors large $\delta$, and defect formation, so that optimality is reached near the regime where defects appear, i.e. for $e^{\delta}$ not too large in comparison with $1/\rho$. However, at that stage structures are still formed, but with a few defects. Thus, if a more relaxed definition of yield that includes structures with a small number of defects is considered, we expect that this optimal relaxed yield is obtained for larger $\delta$. This view is supported by Fig. 12, comparing the yield (left panel) with a relaxed definition of yield, which indeed peaks at a larger $\delta$ (Fig. 12).
@@ IV Conclusion
We have established that for hierarchical self-assembly, various quantities scale with the mass $m$ of sub-structures considered. On the one hand, pairwise interaction should decrease at the boundary of large units, such that structures at a given scale display an overall interaction independent of that scale. For compact sub-structures that fill up space, this implies a microscopic interaction strength decreasing as $m^{(d-1)/d}$ where $d$ is the spatial dimension. We argued that the optimal interaction strength $\delta$ will satisfy $C_{1}/\rho\leq e^{\delta}\leq C_{2}/\rho^{2}$, the first inequality marking where defects starts to increase, and the second where hierarchical assembly breaks down as generations do not grow sequentially. Determining quantitatively the precise location where yield is optimal for a given assembly duration remains a theoretical challenge for the future. Overall, although our predictions are verified in a simple model of sticky squares, we expect them to apply in more general situations, including with DNA bricks where they could be tested.
Our analysis also suggests to revisit the role of hierarchical assembly in protein folding. Folding is believed to occurs successfully because the energy landscape presents a funnel shape Bryngelson et al. (1995), a description that however does not distinguish between very distinct folding mechanisms. Multiple mechanistic views exist, ranging from pure nucleation Dokholyan et al. (2000) to hierarchical folding Baldwin and Rose (1999); Dill et al. (2008) as well as intermediary scenarios including both phenomena acting at different scales Fersht (1997). Studying this question would require to extend our analysis to the case where monomers come from a finite set, and are constrained to form a chain. Interestingly, this problem may become experimentally accessible in soft matter systems, for which chains of particles with specific interactions can be built McMullen et al. (2022).
@@ Acknowledgements.
Appendix A Equilibrium calculations, details
@@
First suppose we wish to make a 4-square, and consider all the intermediate fragment types: there are 4 types of monomers with energy 0, 4 types of 2-mers with energy $-\delta$, 4 types of trimers with energy $-2\delta$, and 1 type of 4-mer with energy $-4\delta$. Let $n_{i}$ be the total number of $i$-mers observed, and let $c_{i}$ be the concentration of each type of $i$-mer, so that $c_{1}=n_{1}/4V$, $c_{2}=n_{2}/4V$, $c_{3}=n_{3}/4V$, $c_{4}=n_{4}/V$. The different species interact with reactions
[…]
@@
$c_{1}=e^{-\delta}c_{4}^{1/4},\qquad c_{2}=e^{-\delta}c_{4}^{1/2},\qquad c_{3}=e^{-\delta}c_{4}^{3/4}.$ (10)
@@
Now suppose we wish to form a $4^{k}$-square using hierarchical interactions. The system must first form 4-squares; therefore we must similarly ask which value of $\delta=\delta_{1}$ makes these 4-squares have high yield, before we turn on the hierarchical interactions (i.e. assuming $\delta_{2}=\delta_{3}=\cdots=0$). The difference from a system of pure 4-squares is that now there are $4^{k-1}$ distinct copies of each 4-square, so the equation for mass conservation (11) is replaced by
$4c_{1}+8c_{2}+12c_{3}+4c_{4}=\frac{\phi}{4^{k-1}}$ (13)
because the overall concentration of each type of monomer is reduced by a factor of 4. Therefore, the concentration $c_{4}=n_{4}/4^{k-1}V$ satisfies (12), but with $\phi\to\phi/4^{k-1}$.
When the yield $y_{1}^{(k)}=4^{k}c_{4}/\phi$ is high, we may obtain a perturbative solution to (13) by letting $y_{1}^{(k)}=1-\Delta_{4}^{(k)}$ with $\Delta_{4}^{(k)}\ll 1$. At low density where $c_{4}\ll 1$, we must have $c_{1}\gg c_{2}\gg c_{3}$, so we drop $c_{2},c_{3}$. Because of the high yield assumption, $c_{4}\gg c_{1}$, so we only include the perturbation in $c_{4}$, making the approximations $c_{4}\approx(1-\Delta_{4}^{(k)})\phi/4^{k}$, $c_{1}\approx(\phi/4^{k})^{1/4}$. Solving (13) using these approximations gives
$\Delta_{4}^{(k)}\approx e^{-\delta}\left(\frac{\phi}{4^{k}}\right)^{-3/4}.$ (14)
@@
$\delta\approx k\frac{3}{4}\log 4-\frac{3}{4}\log\phi-\log\Delta_{4}^{(k)}.$ (15)
Appendix B Kinetics of forming a 4-square
We may solve for the time-dependent concentrations of 4-squares by solving nonlinear ODEs corresponding to the different sub-reactions involved in forming a 4-square. We consider the reactions (9), and assume (i) infinite volume at fixed volume fraction (ii) the system is well-mixed, and (iii) the initial concentrations of each sub-type of fragment are the same (monomer, dimer, trimer). Then standard methods give the system of reaction equations
@@ Appendix B Kinetics of forming a 4-square
We solved (16) numerically to find $\delta_{\rm opt}(t)$, the interaction parameter $\delta_{1}$ that maximizes the concentration $c_{4}(t)$ at fixed $t$, and the corresponding value of the optimal yield, $y_{4}^{\rm opt}(t)=c_{4}(t)/\rho$, at fixed volume fraction $\phi=0.05$. We performed the computations in Julia using packages DifferentialEquations, SciMLSensitivity, OptimJL, to solve the equations numerically, compute the gradient of $c_{4}(t)$ using automatic differentiation, and find the optimal parameter values using the BFGS algorithm. Fig. 3 shows a clear scaling of the optimal energy $e\delta_{\rm opt}(t)\sim t^{0.91}$, and of the optimal yield $1-y_{4}^{\rm opt}(t)\propto t^{-0.97}$. These scalings are consistent with our predictions from (2),(3).
Appendix C Additional figures: additional data, and finite-size effects
@@
This section collects additional figures to support our claims in the main text.
C.1 A system of $n=1024$-squares assembling
@@
Figure 13 shows snapshots of a system of 1024-squares assembling. Computational limitations mean we can only consider 2 copies of this system, however it is notable that we achieve very high assembly yields.
@@
Figure 13: Snapshots of two copies of a 1024-square assembling, using hierarchical interactions with $\delta_{1}=10$.
C.2 Finite size effects
@@
Figure 14: Testing finite size effects in simulations of a 64-square with hierarchical interactions, at various values of $\delta_{1}$. The simulations consider $N_{c}$ copies of a 64-square, with values of $N_{c}$ shown in the legend, and with statistics averaged over $N_{r}$ independent realizations chosen such that there are a total of $N_{c}N_{r}=128$ total copies simulated for each choice of parameters.

### v2

@@ II.4 Assembling $n$-squares with high yields via hierarchical interactions
As a check that we can extend hierarchical assembly by one more generation we simulated $N_{c}=2$ copies of a 1024-square using $\delta_{1}{=}10$, for a time of $1.15\times 10^{6}$ MC sweeps, and repeated 4 times. Out of the 8 total copies of a 1024-square simulated, 7 copies formed perfectly, and 1 copy was missing two monomers from its interior. Snapshots of a perfectly-assembling system are shown in the SI, Fig. S3.1.
Hierarchical assembly as the number of generations $k$ varies is compared in Fig.3(c). To choose the duration of these runs systematically, we consider $k=1$ and determine a time $t_{1}=64\times 10^{3}$ MC sweeps at which $y_{1}$ is close enough to equilibrium over a range of $\delta_{1}$ values. For larger $k$, we chose $t_{k}=8^{k-1}t_{1}$, based on scaling time with the diffusion timescale for each generation, $\tau_{D}^{k}=(D_{k}\rho)^{-1}$ (for each subsequent generation, $\rho$ decreases by a factor of 4 for fixed volume fraction $\phi$, and $D_{k}$ decreases by a factor of 2). Generation $k=5$ is run for about half that time because of computational limitations.
Fig. 3(c) shows that the yields are high over a range of $\delta_{1}$-values for all generations, though they decrease slightly with each new generation as explained below. An exception is $k=5$ which has abnormally high yields, which we attribute to finite-size effects (Fig. S3; SI Section S3.3). The curves appear shifted to the right by about $1k_{B}T$ for each additional generation. This is consistent with (1), which shows that for fixed error rate $\Delta_{1}$ in the first generation and fixed volume fraction $\phi=4^{k}\rho$, the interaction energy must increase by about $3/4\log 4\approx 1.04k_{B}T$ for each additional generation (see also SI, Eq.(S1.7)).
@@ II.5 Dynamics of hierarchical assembly
To quantify this behaviour further, in the left panel of Fig.S2 the mass in fragments of size $n=4,16,64,256$ is shown as a function of time, for a system of assembling 256-squares with optimal hierarchical interactions. The peaks in the masses occur at subsequently increasing times and then decay, except for the targeted structure $n=256$. Noticeably, the decaying parts of the curves appear simply shifted in time by $1/\sqrt{n}$ as $n$ increases (right panel), consistent with scaling the decay with the diffusion timescales as we rationalize below.
@@ III.1 Themodynamics constraints
We seek to estimate the yield $y_{k}$ of hierarchical self-assembly with $k$ generations, as a function of the monomer density $\rho$ and of the interaction strengths. Let us denote by $\tilde{\delta}_{j}=2^{j}\delta_{j}$ the effective interaction between two fully formed squares at the $j^{th}$ intermediate generation, each consisting of $4^{j}$ monomers. We will assume that $\tilde{\delta}_{j}=\delta$ independently of $j$. We denote by $y_{j}^{(k)}$ the yield of such squares. At the first generation, assuming that weaker interactions do not disturb the kinetic of the assembly process (see below), from Eq.1 we get $y_{1}^{(k)}=1-\Delta_{1}(\rho)$. We can now proceed recursively, assuming that (i) the formation of squares at generation $j$ is not affected by that of larger structures and (ii) that it only depends on the concentration of j1j-1 squares obtained in the process, which only depends on the yields of previous stages $j^{\prime}$ with $1\leq j^{\prime}\leq j$. The yields of intermediate generations are thus approximated as
@@
and the target yield $y_{k}=y_{k}^{(k)}$ is computed recursively. This yield approximation is tested against simulation data in Fig.3(c) (dashed lines). It correctly captures the trend for the yield, and even agrees quantitatively for weak interactions and small $k$.
@@ III.2 Kinetic constraints
Remaining hierarchical: For hierarchical assembly to proceed, interactions needed to build the generation $j+1$ should not interfere with the assembly of generation $j$; they should start to play a role only when the $j$-squares are fully formed. Incomplete substructures (such as triangles) of two adjacent $j$-squares bind with an interaction $\delta/2$. They will thus stick together even when the $j$ squares are not completed, and remain attached for some time $\sim e^{\delta/2}/D_{j}$. If this time is much larger than the diffusion time $\tau_{D}^{j}\sim 1/(\rho D_{j})$ for $j$-squares to meet, then partially formed structures spend most of their time bound to other partially formed structures. Proper hierarchical assembly can only happen when this does not occur, implying:
@@
Avoiding defects: As discussed for the formation of 4-squares, the concentration of triangles will be abundant at any level $j$ at some point during the hierarchical assembly if the ratio of unbinding time over diffusion time $f^{j}\equiv\tau^{j}_{B}/\tau^{j}_{D}=e^{\delta}\rho$ is large. If the associated condition $f^{j}\ll 1$ or equivalently:
$e^{\delta}\ll\frac{1}{\rho}$ (8)
does not hold, we expect to see defects composed of a triangle binding to three completed squares, such that an inaccessible hole is present in the interior of this structure. Indeed, the energy of removing the triangle to heal such a defect is precisely $\delta$; condition Eq.8 ensures that the healing time is smaller than the characteristic time where intermediate structures meet and bind.
Interestingly, if we denote by $t_{j}$ the time scale to form generation $j$, we have that $t_{j}$ is always inversely proportional to $D_{j}$. Indeed, $t_{j}$ corresponds to the slowest time between the unbinding time $\tau^{j}_{B}$ and the diffusion time $\tau^{j}$, and both satisfy this property. Since $D_{j}\sim 1/\sqrt{4^{j}}$, one expects that during the hierarchical assembly, subsequent generations should take 2 times longer to form. Thus, $t_{j}\sim 2^{j}\sim\sqrt{n}$, as confirmed in Fig.S2.
Figure 6: Toward a phase diagram for hierarchical self-assembly. (a): A schematic phase diagram, showing regions with different qualitative behaviour. The regions are demarcated by lines of the form (for increasing $\delta)$ (5), (8), (7),(6). (b,c): Contour plots of the yield for $n=64$ with hierarchical interactions as a function of strongest interaction energy $\delta$ and volume fraction $\phi$. The yield includes structures with 0 defects (grey shades / solid lines) and 2 defects (blue shades / dashed lines), at times (b) $t=4.096\times 10^{6}$ and (c) $t=8.192\times 10^{6}$ MC sweeps. Purple dotted lines show $\log(CtD)$ with $C_{1}=1/9$, $C_{2}=(1/35)/64$, and $D=1/16$ the diffusivity for a 16-square. Green dotted line has slope $0.75$ and is a guide to the eye. (d): Snapshots showing qualitatively how hierarchical assembly behaves as $\delta$ increases, for $n=64$ at fixed volume fraction $\phi=0.05$, at time $t=8.192\times 10^{6}$ MC sweeps. Each panel is a portion of one simulation described in Fig. S4with $\phi=0.05$.
@@ III.3 Phase diagram
Our results are summarized in the qualitative phase diagram of Fig. 6 (a). For fixed $\rho$, at low $\delta$ the yield is poor as expected thermodynamically, and improves as $\delta$ increases. The yield becomes large when $e^{\delta}>C/\rho^{3/4}$ as in Eq.(5) (here $C$ represents an unknown constant). For $e^{\delta}>C/\rho$ as in Eq.(8), defects can start to appear. For even stronger interactions $e^{\delta}>C/\rho^{2}$ as in Eq.(7), the notion of hierarchical assembly is lost, as interactions associated with different levels can bind concomitantly. Finally for $e^{\delta}>C\,tD$ as in Eq.(6), low-level bonds have no time to detach: kinetic traps appear already at small scales. Snapshots from different regions of the diagram showing these different qualitative behaviours are shown in Fig. 6(d).
@@
Overall, from these arguments we expect that yield is governed by a trade-off between thermodynamics, which favors large $\delta$, and defect formation, so that optimality is reached near the regime where defects appear, i.e. for $e^{\delta}$ not too large in comparison with $1/\rho$. However, at that stage structures are still formed, but with a few defects. Thus, if a more relaxed definition of yield that includes structures with a small number of defects is considered, we expect that this optimal relaxed yield is obtained for larger $\delta$.
To test these predictions, we simulate systems of $n=64$-squares on intermediate time scales to obtain an empirical phase diagram, as described in more detail in the SI (Section S3.4). Fig. 6 (b,c) show contour plots of the yield (grey shades) as a function of $(-\log\phi,\delta)$ (recall $\phi=64\rho$), for two assembly times separated by a factor of 2. We superimposed contour lines for a relaxed definition of yield (blue shades), which includes structures with at most 2 defects (hence which are at least 97% complete). Key observations include:
  • As predicted, the high-yield region disappears when $\delta$ is too small or too large, and when the volume fraction $\phi$ is too small.
  • Both the threshold inverse packing fraction $1/\phi_{c}$ and the threshold $\delta_{c}$ beyond which the yield degrades are consistent with our predictions $\log(CtD)$, where $C$ is a constant that depends on the quantity considered. This is apparent when comparing panels (b) and (c) of Fig. 6. Specifically, we determine from the middle plot (b) the constants $C_{1},C_{2}$ such that $\delta_{c}=\log(C_{1}tD)$ and $\log(1/\phi_{c})=\log(C_{2}tD)$, where $\delta_{c}$ and $1/\phi_{c}$ are defined as the largest values of $\delta$ and $1/\phi$ for which 50% yield is reached. We can see that our predictions for $\delta_{c}$ and $1/\phi_{c}$ still hold as the time $t$ is changed in panel (c).
  • As predicted, both definitions of yield initially increase at the same rate with $\delta$. However, the relaxed definition of yield that allows for defect remains high for a wider range of $\delta$ values, validating our prediction that the assembly starts to degrade due to vacancy defects.
  • The slopes of the lines where high yield appears are $\approx 0.6{-}0.7$, close to the prediction of Eq.5 of slope $0.75$. We measure similar slopes for the emergence of defects (corresponding to suppression of the high yield region, while the relaxed definition is still high), close to the prediction of slope 1, Eq.8.
The emergence of mixing (7) corresponds to the decrease of the relaxed definition of yield. Measuring it is difficult, as it requires to be in the regime of large $n$ and $t$ which hard to access in simulations; see the discussion in the SI Section S3.4.
III.4 Role of non-native interactions
The yield is expected to degrade when there are non-native interactions present, also called crosstalk. Crosstalk has been shown to significantly degrade the yield of self-assembly once the average crosstalk energy is above a threshold Murugan et al. (2015b), and some amount of crosstalk is inevitable in physical systems Huntley et al. (2016). For hierarchical assembly, we expect that cross-talk will affect mostly the assembly of the largest structures (i.e. the squares of size $n/4$). If the cross-talk energy is a constant $U$, these structures will bind non-specifically with an energy $\sim U\sqrt{n}$ (neglecting numerical pre-factors), thus the threshold cross-talk energy $U_{c}(n)$ beyond which self-assembly is impossible follows $U_{c}\sim 1/\sqrt{n}$. If the cross-talk energies are of mean zero and standard deviation $\sigma$, then problems are milder and can be estimated with similar arguments. In particular, the threshold for assembly expected to scale as $\sigma_{c}\sim n^{-1/4}$.
To check the effect of crosstalk we performed simulations for different $n$-squares with different levels of crosstalk, assumed to be of constant magnitude (details in SI Section S3.5). For each $n$, the yield was unaffected for crosstalk energy below a threshold $U_{c}(n)$, beyond which it decreased rapidly (Fig.S6). The threshhold follows approximately our prediction $n^{-1/2}$.
@@ IV Conclusion
We have established that for hierarchical self-assembly, various quantities scale with the mass $m$ of sub-structures considered. On the one hand, pairwise interaction should decrease at the boundary of large units, such that structures at a given scale display an overall interaction independent of that scale. For compact sub-structures that fill up space, this implies a microscopic interaction strength decreasing as $m^{(d-1)/d}$ where $d$ is the spatial dimension. We argued that the optimal interaction strength $\delta$ will lie in some interval, with a lower bound given by thermodynamic considerations, and an upper bound determined when the diffusion time becomes slower than the unbinding time and hence vacancy defects start to appear. Determining quantitatively the precise location where yield is optimal for a given assembly duration remains a theoretical challenge for the future.
Overall, we have shown that hierarchical assembly is a highly successful strategy in a simple model of sticky squares, but we expect it to be a useful principle in more general situations. Future work can explore how to extend these ideas to systems with different geometries, including to DNA bricks where our predictions could be tested experimentally.
Our analysis also suggests to revisit the role of hierarchical assembly in protein folding. Folding is believed to occurs successfully because the energy landscape presents a funnel shape Bryngelson et al. (1995), a description that however does not distinguish between very distinct folding mechanisms. Multiple mechanistic views exist, ranging from pure nucleation Dokholyan et al. (2000) to hierarchical folding Baldwin and Rose (1999); Dill et al. (2008) as well as intermediary scenarios including both phenomena acting at different scales Fersht (1997). Studying this question would require extending our analysis to the case where monomers come from a finite set, and are constrained to form a chain. Interestingly, this problem may become experimentally accessible in soft matter systems, for which chains of particles with specific interactions can be built McMullen et al. (2022).
@@ Acknowledgements.
Appendix S1 Equilibrium calculations, details
@@
First suppose we wish to make a 4-square, and consider all the intermediate fragment types: there are 4 types of monomers with energy 0, 4 types of 2-mers with energy $\delta$, 4 types of trimers with energy $2\delta$, and 1 type of 4-mer with energy $4\delta$. Let $n_{i}$ be the total number of $i$-mers observed, and let $c_{i}$ be the concentration of each type of $i$-mer, so that $c_{1}=n_{1}/4V$, $c_{2}=n_{2}/4V$, $c_{3}=n_{3}/4V$, $c_{4}=n_{4}/V$. The different species interact with reactions
$\text{1-mer}+\text{1-mer}$ $\leftrightarrow\text{2-mer},$
$\text{1-mer}+\text{2-mer}$ $\leftrightarrow\text{3-mer}$
$\text{2-mer}+\text{2-mer}$ $\leftrightarrow\text{4-mer},$
$\text{1-mer}+\text{3-mer}$ $\leftrightarrow\text{4-mer}.$ (S1.1)
@@
$c_{1}=e^{-\delta}c_{4}^{1/4},\qquad c_{2}=e^{-\delta}c_{4}^{1/2},\qquad c_{3}=e^{-\delta}c_{4}^{3/4}.$ (S1.2)
@@
Now suppose we wish to form a $4^{k}$-square using hierarchical interactions. The system must first form 4-squares; therefore we must similarly ask which value of $\delta=\delta_{1}$ makes these 4-squares have high yield, before we turn on the hierarchical interactions (i.e. assuming $\delta_{2}=\delta_{3}=\cdots=0$). The difference from a system of pure 4-squares is that now there are $4^{k-1}$ distinct copies of each 4-square, so the equation for mass conservation (S1.3) is replaced by
$4c_{1}+8c_{2}+12c_{3}+4c_{4}=\frac{\phi}{4^{k-1}}$ (S1.5)
because the overall concentration of each type of monomer is reduced by a factor of 4. Therefore, the concentration $c_{4}=n_{4}/4^{k-1}V$ satisfies (S1.4), but with $\phi\to\phi/4^{k-1}$.
When the yield $y_{1}^{(k)}=4^{k}c_{4}/\phi$ is high, we may obtain a perturbative solution to (S1.5) by letting $y_{1}^{(k)}=1-\Delta_{4}^{(k)}$ with $\Delta_{4}^{(k)}\ll 1$. At low density where $c_{4}\ll 1$, we must have $c_{1}\gg c_{2}\gg c_{3}$, so we drop $c_{2},c_{3}$. Because of the high yield assumption, $c_{4}\gg c_{1}$, so we only include the perturbation in $c_{4}$, making the approximations $c_{4}\approx(1-\Delta_{4}^{(k)})\phi/4^{k}$, $c_{1}\approx(\phi/4^{k})^{1/4}$. Solving (S1.5) using these approximations gives
$\Delta_{4}^{(k)}\approx e^{-\delta}\left(\frac{\phi}{4^{k}}\right)^{-3/4}.$ (S1.6)
@@
$\delta\approx k\frac{3}{4}\log 4-\frac{3}{4}\log\phi-\log\Delta_{4}^{(k)}.$ (S1.7)
Appendix S2 Kinetics of forming a 4-square
We may solve for the time-dependent concentrations of 4-squares by solving nonlinear ODEs corresponding to the different sub-reactions involved in forming a 4-square. We consider the reactions (S1.1), and assume (i) infinite volume at fixed volume fraction (ii) the system is well-mixed, and (iii) the initial concentrations of each sub-type of fragment are the same (monomer, dimer, trimer). Then standard methods give the system of reaction equations
@@
We solved (S2.1) numerically to find $\delta_{\rm opt}(t)$, the interaction parameter $\delta_{1}$ that maximizes the concentration $c_{4}(t)$ at fixed $t$, and the corresponding value of the optimal yield, $y_{4}^{\rm opt}(t)=c_{4}(t)/\rho$, at fixed volume fraction $\phi=0.05$. We performed the computations in Julia using packages DifferentialEquations, SciMLSensitivity, OptimJL, to solve the equations numerically, compute the gradient of $c_{4}(t)$ using automatic differentiation, and find the optimal parameter values using the BFGS algorithm. Fig. 2 shows a clear scaling of the optimal energy $\delta_{\rm opt}(t)\sim t^{0.91}$, and of the optimal yield $1-y_{4}^{\rm opt}(t)\propto t^{-0.97}$. These scalings are consistent with our predictions from (2),(3).
Appendix S3 Additional figures and data
@@
This section collects additional figures and data to support our claims in the main text.
S3.1 A system of $n=1024$-squares assembling
@@
Figure S3.1 shows snapshots of a system of 1024-squares assembling with hierarchical interactions. Computational limitations mean we can only consider 2 copies of this system, however it is notable that we achieve very high assembly yields.
@@
Figure S3.1: Snapshots of $N_{c}=2$ copies of a 1024-square assembling, using hierarchical interactions with $\delta_{1}=10$. Each panel shows the full simulation.
S3.2 Mass versus time
Figure S3.2: (Left) Fraction of the system mass in fragments of size $n=4,16,64,256$, as a function of time, for a system of $N_{c}{=}16$ 256-squares with $\delta_{1}=8.75$, repeated 8 times. Time is measured in units of $10^{4}$ MC sweeps. (Right) Same data but with time rescaled by $1/\sqrt{n}$.
Figure S3.2 shows the fraction of the system’s mass in fragments of size $n=4,16,64,256$, as a function of time, for a system of 256-squares. The first plot shows the raw data, while the second rescales time by $t\to t/\sqrt{n}$. The decay rates align for this rescaling, supporting our claim that subsequent generations in this hierarchical assembly process should take 2 times longer to form.
S3.3 Finite size effects
@@
Figure S3.3: Testing finite size effects in simulations of a 64-square with hierarchical interactions, at various values of $\delta_{1}$. The simulations consider $N_{c}$ copies of a 64-square, with values of $N_{c}$ shown in the legend, and with statistics averaged over $N_{r}$ independent realizations chosen such that there are a total of $N_{c}N_{r}=128$ total copies simulated for each choice of parameters. Vertical lines are one-standard deviation error bars.